OpenStudy (anonymous):

The rational function \(c(x) := \frac{x+1}{x-1}\) is called Cayley transformation d) Is \(c^{-1}\) continuous?

OpenStudy (blockcolder):

What definition of continuity do you know?

OpenStudy (anonymous):

a continuous function is a function for which, intuitively, "small" changes in the input result in "small" changes in the output. Otherwise, a function is said to be "discontinuous". A continuous function with a continuous inverse function is called "bicontinuous".

OpenStudy (anonymous):

is it?

OpenStudy (blockcolder):

There are many ways to prove continuity of a function. Here are 3 of them: 1. Prove that \(\lim_{x \to a}f(x)=f(a)\). 2. Prove that \(f'(a)\) exists for all a in the domain. 3. Prove that \(\lim_{h \to 0}f(a+h)=f(a)\).

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

wich is the better one in our situation?

OpenStudy (blockcolder):

Try Method 2. It is my preferred one, actually.

OpenStudy (anonymous):

blockcolder the problem is i dont have enough time to solve it i will be very happy if you can do it for me i must give it to mentor in 2-3 hours

OpenStudy (blockcolder):

What's the inverse function of c(x)? Can you write it here?

OpenStudy (anonymous):

\[c(x)^{-1}\] ?

OpenStudy (blockcolder):

Yeah. Write its equation here.

OpenStudy (anonymous):

\[c(x)^{-1} =c \frac{1}{x} = \]

OpenStudy (anonymous):

i am stucking now..

OpenStudy (blockcolder):

Did you understand what I said in your previous question?

OpenStudy (anonymous):

hmm, i thought it was complete solution, do i need to write something additional??

OpenStudy (blockcolder):

I thought you worked it out on your own. What I said back then was just how to do it.

OpenStudy (anonymous):

no unfortunately for mathematics have really less time..but in one and half month i have exam i must start to learn good analysis math

OpenStudy (anonymous):

but i am allowed to write math exam just if i have enough pooints..

OpenStudy (blockcolder):

You're lucky; the function is its own inverse, as I just discovered a while ago. So now what you have to do is prove that \(c^{-1}(x)=\frac{x+1}{x-1}\) is continuous. For this one, method 1 gives the fastest solution. \[\begin{align}\lim_{x \to a}f(x)&=\lim_{x \to a} \frac{x+1}{x-1}\\ &=\lim_{x \to a}(x+1) \cdot \lim_{x \to a}\frac{1}{x-1}\\ &=\lim_{x \to a}(x+1) \cdot \frac{1}{\lim_{x \to a}x-1}\\ &=(a+1)\cdot \frac{1}{a-1}\\ &=\frac{a+1}{a-1}\\ &=f(a) \end{align}\] Hopefully you know the limit rules I used here?

OpenStudy (anonymous):

to be honest i dont know very well but in first step i need to points for exam participation thank you very much